When the palmaris longus muscle in the forearm is flexed, the wrist moves back and forth. If the muscle generates a force of 39.5 N and it is acting with an effective lever arm of 2.55 cm, what is the torque that the muscle produces on the wrist?

When the palmaris longus muscle in the forearm is

When the palmaris longus muscle in the forearm is
Answer #1

Torque produced by muscle is,

$$ \begin{array}{l} \tau=r_{\text {eff }} F \\ \tau=(2.55 \mathrm{~cm})(39.5 \mathrm{~N}) \\ \tau=\left(2.55 \times 10^{-2} \mathrm{~m}\right)(39.5 \mathrm{~N}) \\ \tau=1.01 \mathrm{~N} \cdot \mathrm{m} \end{array} $$